# Switch-JustDance: Benchmarking Whole-Body Motion Tracking Policies Using a Commercial Console Game

### v2

@@ 1 Introduction
Recent progress in whole-body robot control has enabled humanoid and legged robots to demonstrate increasingly agile, coordinated, and human-like movements. Advances in both model-based whole-body controllers [khazoom2024tailoring, li2024cafe] and reinforcement learning [xie2025kungfubot, chen2025gmt, li2025amo, ze2025twist] have led to robots executing skills such as gymnastic motions, dynamic dancing, and real-time teleoperation. These developments mark a significant step toward embodied agents that can interact robustly with the world in athletic and versatile ways. As these capabilities evolve, a new and fundamental challenge is emerging for the community: how to evaluate the athletic performance of robot control policies in a fair, reproducible, and physically grounded way that allows direct comparison with humans.
@@
Many recent works rely on large-scale human motion datasets such as AMASS [AMASS:ICCV:2019] and LaFAN1 [harvey2020robust] to train motion imitation policies and evaluate whole-body controllers for humanoid and legged robots [he2024hover, Sferrazza2024HumanoidBench, Tessler2024MaskedMimic, shrestha2024generating]. While these datasets have been critical for enabling diverse motion learning, they are infrequently updated, and the choice of splits and evaluation protocols varies across studies, making consistent cross-method comparison difficult. Simulation benchmarks such as HumanoidBench [Sferrazza2024HumanoidBench] begin to standardize tasks, but remain restricted to specific robot models and simulated environmen ts. Real-world benchmarking is even more challenging: motion-capture-based evaluation requires dedicated equipment, controlled environments, and high operational cost, and therefore large-scale or community-wide adoption remains impractical. Finally, robotic performance is rarely compared directly against humans under the same evaluation criteria, leaving open the question of how close robots are to approaching human-level athletic skill. These issues highlight the need for a standardized, reproducible, and accessible framework for benchmarking whole-body control in real-world conditions.
@@
We present Switch-JustDance, a benchmarking framework that leverages the commercial motion-sensing game Just Dance on the Nintendo Switch [nintendoSwitch, justdance] to evaluate whole-body humanoid controllers. The Nintendo Switch provides a low-cost ($\sim\mathdollar 400$), plug-and-play setup that is robust for repeated use, while Just Dance offers a large and continually updated library of full-body motions and a built-in scoring system that naturally accommodates players of different body sizes. Because the platform is designed for human users, it enables direct and quantitative comparisons between robots and humans under identical conditions, making it a promising candidate for standardized evaluation. In this work, we take a first step toward establishing its benchmarking potential by conducting a systematic validation study of the Just Dance scoring system. We analyze its validity, discriminability, and repeatability through an extensive user study. Our results demonstrate that the scoring mechanism, which relies on the IMU-based Joy-Con controller, produces stable and interpretable performance measures, supporting its use as a reliable benchmark for embodied AI.
@@
Building on this foundation, we instantiate Switch-JustDance on a humanoid platform and benchmark three representative state-of-the-art controllers: GMT [chen2025gmt], TWIST [ze2025twist], and Any2Track [zhang2025any2track]. By evaluating the performance of these controllers and comparing them against human performance, we show that current humanoid controllers still fall far short of human-level ability, and that Switch-JustDance reveals several important performance characteristics: smoothed input motions improve tracking stability, dynamic motions expose the limits of high-frequency whole-body control, and simulation results tend to overestimate controller robustness compared to hardware. These findings highlight both the challenges of dynamic whole-body dancing and the value of Switch-JustDance as a fair and physically grounded benchmarking tool.
@@ 2.1 Humanoid whole-body controllers
Recent years have seen rapid progress in learning-based and optimization-based controllers for humanoid whole-body control. Learning approaches include versatile generalist policies that unify multiple control modes via motion-imitation distillation [he2024hover], teleoperation-to-policy pipelines that track human motion on hardware [ze2025twist], unified motion-tracking frameworks trained from large mocap datasets [chen2025gmt], hybrid RL+trajectory-optimization methods for hyper-dexterous movements [li2025amo], and controllers targeting highly dynamic skills such as kung-fu and dancing [xie2025kungfubot]. In parallel, optimization-centric systems push whole-body model predictive control toward real-time deployment, including cascaded-fidelity MPC with tuning-free whole-body control [li2024cafe] and accuracy-tailored fast whole-body NMPC for legged/humanoid robots [khazoom2024tailoring].
Evaluation in these works typically combines (i) simulation metrics—pose/joint/EE tracking errors, stability and energy proxies—using mocap or keyframe references; and (ii) hardware demonstrations scored via task success or qualitative videos, sometimes with additional tracking metrics from external motion capture [he2024hover, chen2025gmt, li2025amo, xie2025kungfubot]. While these practices validate feasibility and absolute tracking quality, they are difficult to compare across papers due to heterogeneous datasets, reward/metric definitions, and task suites. Moreover, real-world evaluations that rely on optical MoCap demand dedicated spaces and instrumentation, which are costly and cumbersome to replicate at scale. As a result, many studies remain partly simulation-bound and rarely provide a human-vs-robot comparison under identical scoring—leaving open the question of how close current controllers are to human athletic performance.
@@ 2.2 Commercial games for AI benchmarking
Commercial games have long catalyzed progress in AI benchmarking and dataset creation. The Arcade Learning Environment (ALE) enabled reproducible, multi-task evaluation across dozens of Atari 2600 games [bellemare2013arcade], underpinning breakthroughs such as DQN’s human-level control [mnih2015human]. Subsequent platforms broadened task complexity and partial observability (e.g., VizDoom [kempka2016vizdoom] and StarCraft II [vinyals2017starcraft]). Minecraft-based ecosystems advance open-ended embodied agents and sample-efficient RL through competitions and large knowledge bases (MineRL and MineDojo) [milani2020retrospective, kanervisto2022minerl, fan2022minedojo]. For driving perception, Grand Theft Auto V has been exploited to generate large, pixel-accurate synthetic datasets and labels at scale, accelerating training and reducing annotation cost [richter2016playing]. Across these examples, games provide standardized interfaces, abundant content, and community baselines. Our work follows this lineage but targets physically grounded, hardware-in-the-loop benchmarking for embodied AI.
@@ 2.3 VR and console platforms for embodied AI
Commodity gaming hardware has long been used to lower barriers in embodied AI research. Microsoft Kinect enabled low-cost full-body motion capture and was widely adopted for human–robot interaction, gesture-based teleoperation, and benchmarking tasks [shotton2011real, sturm2012benchmark]. Similarly, Nintendo Wii devices, including Wii-motes and the Balance Board, were employed in rehabilitation and HRI studies to capture motion and balance signals for interactive training [burke2009optimising]. These works demonstrated that consumer-level consoles can meaningfully support motion capture and embodied evaluation, albeit often with domain-specific or limited reproducibility. More recently, VR-based platforms have emerged as flexible alternatives to optical MoCap for humanoid learning [winkler2022questsim, ahuja2022controllerpose]. Systems such as Open-TeleVision [cheng2024open], H2O [he2024learning], OmniH2O [he2024omnih2o], and OpenWBT [zhang2025openwbt] enable immersive whole-body teleoperation and large-scale demonstration collection for humanoid robots using head-mounted displays and consumer sensors. These pipelines provide naturalistic embodiment and scalable data collection, but still require careful calibration and custom infrastructure. Our work builds on this trajectory by exploring motion-sensing console games with built-in scoring. In particular, we investigate Just Dance on the Nintendo Switch, which uniquely combines continuous motion updates, standardized quantitative evaluation, and direct human–robot comparability—positioning it as a promising benchmark for whole-body robotic control.
@@ 3 Switch-JustDance
  • MoCap Module: Reconstructs 3D human poses from monocular RGB video using GVHMR [GVHMR], which outputs SMPL [SMPL:2015] parameterizations of the dancer’s motion. We modify GVHMR for real-time operation using a sliding-window inference scheme, achieving stable pose estimates at approximately 200 ms per frame.
  • Retargeting Module: Maps the reconstructed SMPL motions to the target robot morphology using GMR [ze2025gmr], ensuring kinematic feasibility across robots of varying proportions. GMR runs at roughly 50 ms per frame and produces robot-space keyframe trajectories synchronized with the MoCap updates. By using an asynchronous interpolation thread, intermediate frames are interpolated using linear interpolation (LERP) for positions and spherical interpolation (SLERP) for rotations, increasing the output frequency between GVHMR/GMR frames.
  • Motion Tracking Policy: Generates control commands based on the retargeted reference motion and the robot’s proprioceptive feedback. This policy represents the humanoid controller under evaluation.
@@ 4 System Validation
To assess the validity of the Just Dance Score (JDS) as a quantitative metric for humanoid motion tracking, we compare the collected JDS values against a standard kinematic-based metric commonly used in motion tracking evaluations: Procrustes-Aligned Mean Per-Joint Position Error (PA-MPJPE)[Kendall1989ShapeSurvey]. PA-MPJPE measures the average joint position error between the player and a reference motion after applying Procrustes transformation (rigid and scale) in the player’s coordinate. This metric is computed by aligning each participant’s recorded motion with the reference choreography reconstructed by the MoCap Module and then calculating the mean joint position error to quantify tracking accuracy.
Our analysis shows that the correlation coefficient ($r$) and significance level ($p$) between JDS and PA-MPJPE across different songs range from r=[0.76,0.42]r=[-0.76,-0.42] and $p=[<0.001,<0.05]$, clearly indicating that higher JDS values correspond to lower tracking errors, validating JDS as a reliable indicator of motion-tracking quality. Interestingly, the results also imply that JDS captures additional aspects of performance beyond pure kinematic accuracy. According to the related patent document [us9358456b1], the JDS computation also considers features such as joint acceleration and rhythmic alignment with the music beat—important perceptual and dynamic qualities not reflected in PA-MPJPE alone.
@@
To assess the repeatability of the JDS, we examined how consistently it scores repeated executions of the same dance motion. We computed the Intraclass Correlation Coefficient (ICC)[shrout1979intraclass] and the Coefficient of Variation (CV)[fisher1970statistical]. A higher ICC (closer to 1.0) and a lower CV indicate stronger consistency in repeated measurements. Our analysis shows that the JDS achieved an ICC of 0.735 and a CV of 10.0%, indicating low trial-to-trial variation and demonstrating that the metric provides reliable, repeatable evaluations.
In addition to ICC and CV, we computed Kendall’s Coefficient of Concordance (KCC)[kendall1939problem] across multiple songs and repetitions to assess ranking consistency among participants. KCC quantifies the agreement in rank ordering across different conditions, where $KCC=1$ indicates perfect agreement and $KCC=0$ represents random ranking. The obtained $KCC=0.678$ demonstrates that participants who ranked higher in one dance tended to maintain similar rankings across other dances and repetitions. Together, these results confirm the strong repeatability and robustness of the Switch-JustDance benchmarking pipeline.
@@
One limitation of the current benchmarking pipeline is that the Just Dance scoring system relies primarily on the motion of a single Joy-Con, allowing players to achieve high scores through upper-body motion while keeping the lower body static. However, we believe that exploiting the system with only upper-body motion will require careful coordination of trajectories to compensate for the absence of full-body movement, whereas simply leveraging full-body motion holds the promise of achieving higher scores. In future scenarios, we encourage participants to provide videos to ensure visual fidelity.
@@ 5.1 Benchmark Metrics
  • Smoothness: Smoothness is quantified by the average joint jerk(rad/s3[jerk] and acceleration(rad/s2) of the joint angles computed using finite difference. Controllers that produce jittery or impulsive motion yield higher jerk, even when MPJPE is low. This metric emphasizes natural, human-like movement and hardware-friendly actuation.
@@ 5.2 Experimental Settings
[…]
Easy Hard All Active All Jerk(rad/s3) Acc(rad/s2)
Human-leaderBoard 13333 13333 13333 - - 100.0 - -
Human 9867 6542 8546 112.7 112.7 100.0 3275.3 92.5
GMT-Real-Smo 2765 1891 2415 130.9 139.5 73.3 3174.1 78.5
TWIST-Real-Smo 2085 1557 1874 117.1 119.8 80.0 2811.3 70.6
Any2Track-Real-Smo 2978 2785 2900 102.6 139.4 60.0 1978.8 82.6
GMT-Real-Dyn 2881 2530 2741 88.8 102.6 80.0 4098.6 126.3
TWIST-Real-Dyn 1501 1140 1357 108.5 142.7 33.3 4024.1 126.0
Any2Track-Real-Dyn 2753 2251 2552 114.2 128.2 80.0 3783.2 127.6
GMT-Sim-Smo - - - 82.8 137.4 20.0 2146.2 58.7
TWIST-Sim-Smo - - - 116.7 125.2 80.0 1684.6 49.4
Any2Track-Sim-Smo - - - 91.4 106.9 80.0 2115.1 88.1
GMT-Sim-Dyn - - - 87.8 137.0 0.0 3694.6 110.5
TWIST-Sim-Dyn - - - 100.9 127.8 40.0 3438.9 110.1
Any2Track-Sim-Dyn - - - 90.9 105.7 80.0 3949.8 130.1
Table 1: Benchmark results across controllers, input types, and simulation settings. We evaluate three representative humanoid controllers (GMT, TWIST, and Any2Track) under smoothed (Smo) and dynamic (Dyn) input motions, as well as their simulation counterparts. Higher Just Dance Score (JDS) indicates better task performance, lower MPJPE indicates more accurate motion tracking, and higher Success Rate (SR) reflects improved stability. Smoothness is measured using acceleration and jerk.

### v3

# Switch-JustDance: Benchmarking Whole-Body Motion Tracking Controllers Using a Commercial Console Game

@@ Abstract
footnotetext: Project page: https://switch-justdance.github.io/
@@ 1 Introduction
Recent progress in whole-body robot control has enabled humanoid and legged robots to demonstrate increasingly agile, coordinated, and human-like movements. Advances in both model-based whole-body controllers [18, 19] and reinforcement learning [37, 5, 20, 39] have led to robots executing skills such as gymnastic motions, dynamic dancing, and real-time teleoperation. These developments mark a significant step toward embodied agents that can interact robustly with the world in athletic and versatile ways. As these capabilities evolve, a new and fundamental challenge is emerging for the community: how to evaluate the athletic performance of robot control policies in a fair, reproducible, and physically grounded way that allows direct comparison with humans.
@@
Many recent works rely on large-scale human motion datasets such as AMASS [22] and LaFAN1 [10] to train motion imitation policies and evaluate whole-body controllers for humanoid and legged robots [13, 27, 33, 30]. While these datasets have been critical for enabling diverse motion learning, they are infrequently updated, and the choice of splits and evaluation protocols varies across studies, making consistent cross-method comparison difficult. Simulation benchmarks such as HumanoidBench [27] begin to standardize tasks, but remain restricted to specific robot models and simulated environmen ts. Real-world benchmarking is even more challenging: motion-capture-based evaluation requires dedicated equipment, controlled environments, and high operational cost, and therefore large-scale or community-wide adoption remains impractical. Finally, robotic performance is rarely compared directly against humans under the same evaluation criteria, leaving open the question of how close robots are to approaching human-level athletic skill. These issues highlight the need for a standardized, reproducible, and accessible framework for benchmarking whole-body control in real-world conditions.
@@
We present Switch-JustDance, a benchmarking framework that leverages the commercial motion-sensing game Just Dance on the Nintendo Switch [25, 34] to evaluate whole-body humanoid controllers. The Nintendo Switch provides a low-cost ($\sim\mathdollar 400$), plug-and-play setup that is robust for repeated use, while Just Dance offers a large and continually updated library of full-body motions and a built-in scoring system that naturally accommodates players of different body sizes. Because the platform is designed for human users, it enables direct and quantitative comparisons between robots and humans under identical conditions, making it a promising candidate for standardized evaluation. In this work, we take a first step toward establishing its benchmarking potential by conducting a systematic validation study of the Just Dance scoring system. We analyze its validity, discriminability, and repeatability through an extensive user study. Our results demonstrate that the scoring mechanism, which relies on the IMU-based Joy-Con controller, produces stable and interpretable performance measures, supporting its use as a reliable benchmark for embodied AI.
@@
Building on this foundation, we instantiate Switch-JustDance on a humanoid platform and benchmark three representative state-of-the-art controllers: GMT [5], TWIST [39], and Any2Track [41]. By evaluating the performance of these controllers and comparing them against human performance, we show that current humanoid controllers still fall far short of human-level ability, and that Switch-JustDance reveals several important performance characteristics: smoothed input motions improve tracking stability, dynamic motions expose the limits of high-frequency whole-body control, and simulation results tend to overestimate controller robustness compared to hardware. These findings highlight both the challenges of dynamic whole-body dancing and the value of Switch-JustDance as a fair and physically grounded benchmarking tool.
@@ 2.1 Humanoid whole-body controllers
Recent years have seen rapid progress in learning-based and optimization-based controllers for humanoid whole-body control. Learning approaches include versatile generalist policies that unify multiple control modes via motion-imitation distillation [13], teleoperation-to-policy pipelines that track human motion on hardware [39], unified motion-tracking frameworks trained from large mocap datasets [5], hybrid RL+trajectory-optimization methods for hyper-dexterous movements [20], and controllers targeting highly dynamic skills such as kung-fu and dancing [37]. In parallel, optimization-centric systems push whole-body model predictive control toward real-time deployment, including cascaded-fidelity MPC with tuning-free whole-body control [19] and accuracy-tailored fast whole-body NMPC for legged/humanoid robots [18].
Evaluation in these works typically combines (i) simulation metrics—pose/joint/EE tracking errors, stability and energy proxies—using mocap or keyframe references; and (ii) hardware demonstrations scored via task success or qualitative videos, sometimes with additional tracking metrics from external motion capture [13, 5, 20, 37]. While these practices validate feasibility and absolute tracking quality, they are difficult to compare across papers due to heterogeneous datasets, reward/metric definitions, and task suites. Moreover, real-world evaluations that rely on optical MoCap demand dedicated spaces and instrumentation, which are costly and cumbersome to replicate at scale. As a result, many studies remain partly simulation-bound and rarely provide a human-vs-robot comparison under identical scoring—leaving open the question of how close current controllers are to human athletic performance.
@@ 2.2 Commercial games for AI benchmarking
Commercial games have long catalyzed progress in AI benchmarking and dataset creation. The Arcade Learning Environment (ALE) enabled reproducible, multi-task evaluation across dozens of Atari 2600 games [2], underpinning breakthroughs such as DQN’s human-level control [24]. Subsequent platforms broadened task complexity and partial observability (e.g., VizDoom [15] and StarCraft II [35]). Minecraft-based ecosystems advance open-ended embodied agents and sample-efficient RL through competitions and large knowledge bases (MineRL and MineDojo) [23, 14, 7]. For driving perception, Grand Theft Auto V has been exploited to generate large, pixel-accurate synthetic datasets and labels at scale, accelerating training and reducing annotation cost [26]. Across these examples, games provide standardized interfaces, abundant content, and community baselines. Our work follows this lineage but targets physically grounded, hardware-in-the-loop benchmarking for embodied AI.
@@ 2.3 VR and console platforms for embodied AI
Commodity gaming hardware has long been used to lower barriers in embodied AI research. Microsoft Kinect enabled low-cost full-body motion capture and was widely adopted for human–robot interaction, gesture-based teleoperation, and benchmarking tasks [29, 32]. Similarly, Nintendo Wii devices, including Wii-motes and the Balance Board, were employed in rehabilitation and HRI studies to capture motion and balance signals for interactive training [3]. These works demonstrated that consumer-level consoles can meaningfully support motion capture and embodied evaluation, albeit often with domain-specific or limited reproducibility. More recently, VR-based platforms have emerged as flexible alternatives to optical MoCap for humanoid learning [36, 1]. Systems such as Open-TeleVision [6], H2O [12], OmniH2O [11], and OpenWBT [40] enable immersive whole-body teleoperation and large-scale demonstration collection for humanoid robots using head-mounted displays and consumer sensors. These pipelines provide naturalistic embodiment and scalable data collection, but still require careful calibration and custom infrastructure. Our work builds on this trajectory by exploring motion-sensing console games with built-in scoring. In particular, we investigate Just Dance on the Nintendo Switch, which uniquely combines continuous motion updates, standardized quantitative evaluation, and direct human–robot comparability—positioning it as a promising benchmark for whole-body robotic control.
@@ 3 Switch-JustDance
  • MoCap Module: Reconstructs 3D human poses from monocular RGB video using GVHMR [28], which outputs SMPL [21] parameterizations of the dancer’s motion. We modify GVHMR for real-time operation using a sliding-window inference scheme, achieving stable pose estimates at approximately 200 ms per frame.
  • Retargeting Module: Maps the reconstructed SMPL motions to the target robot morphology using GMR [38], ensuring kinematic feasibility across robots of varying proportions. GMR runs at roughly 50 ms per frame and produces robot-space keyframe trajectories synchronized with the MoCap updates. By using an asynchronous interpolation thread, intermediate frames are interpolated using linear interpolation (LERP) for positions and spherical interpolation (SLERP) for rotations, increasing the output frequency between GVHMR/GMR frames.
  • Motion Tracking Controller: Generates control commands based on the retargeted reference motion and the robot’s proprioceptive feedback. This control policy represents the humanoid controller under evaluation.
@@ 4 System Validation
To assess the validity of the Just Dance Score (JDS) as a quantitative metric for humanoid motion tracking, we compare the collected JDS values against a standard kinematic-based metric commonly used in motion tracking evaluations: Procrustes-Aligned Mean Per-Joint Position Error (PA-MPJPE)[16]. PA-MPJPE measures the average joint position error between the player and a reference motion after applying Procrustes transformation (rigid and scale) in the player’s coordinate. This metric is computed by aligning each participant’s recorded motion with the reference choreography reconstructed by the MoCap Module and then calculating the mean joint position error to quantify tracking accuracy.
Our analysis shows that the correlation coefficient ($r$) and significance level ($p$) between JDS and PA-MPJPE across different songs range from r=[0.76,0.42]r=[-0.76,-0.42] and $p=[<0.001,<0.05]$, clearly indicating that higher JDS values correspond to lower tracking errors, validating JDS as a reliable indicator of motion-tracking quality. Interestingly, the results also imply that JDS captures additional aspects of performance beyond pure kinematic accuracy. According to the related patent document [4], the JDS computation also considers features such as joint acceleration and rhythmic alignment with the music beat—important perceptual and dynamic qualities not reflected in PA-MPJPE alone.
@@
To assess the repeatability of the JDS, we examined how consistently it scores repeated executions of the same dance motion. We computed the Intraclass Correlation Coefficient (ICC)[31] and the Coefficient of Variation (CV)[8]. A higher ICC (closer to 1.0) and a lower CV indicate stronger consistency in repeated measurements. Our analysis shows that the JDS achieved an ICC of 0.735 and a CV of 10.0%, indicating low trial-to-trial variation and demonstrating that the metric provides reliable, repeatable evaluations.
In addition to ICC and CV, we computed Kendall’s Coefficient of Concordance (KCC)[17] across multiple songs and repetitions to assess ranking consistency among participants. KCC quantifies the agreement in rank ordering across different conditions, where $KCC=1$ indicates perfect agreement and $KCC=0$ represents random ranking. The obtained $KCC=0.678$ demonstrates that participants who ranked higher in one dance tended to maintain similar rankings across other dances and repetitions. Together, these results confirm the strong repeatability and robustness of the Switch-JustDance benchmarking pipeline.
@@
One limitation of the current benchmarking pipeline is that the Just Dance scoring system relies primarily on the motion of a single Joy-Con, allowing players to achieve high scores through upper-body motion while keeping the lower body static. However, achieving high scores implicitly requires coherent full-body coordination, as rhythm, balance, and timing in dance are strongly coupled under the assumption that the control policy is tasked to faithfully track the reference motion. In future scenarios, we encourage participants to provide videos to ensure visual fidelity.
@@ 5.1 Benchmark Metrics
  • Smoothness: Smoothness is quantified by the average joint jerk(rad/s3[9] and acceleration(rad/s2) of the joint angles computed using finite difference. Controllers that produce jittery or impulsive motion yield higher jerk, even when MPJPE is low. This metric emphasizes natural, human-like movement and hardware-friendly actuation.
@@ 5.2 Experimental Settings
Player-Setting JDS $\uparrow$ MPJPE($m$) $\downarrow$ DTW($m$) $\downarrow$ SR(%) $\uparrow$ Smoothness $\downarrow$
Easy Hard All Active All Act All Jerk(rad/s3) Acc(rad/s2)
Human-leaderBoard 13333 13333 13333 - - - - 100.0 - -
Human 9867 6543 8538 111.9 111.9 102.4 102.4 100.0 3350.5 95.1
GMT-Real-Smo 2765 1891 2415 130.9 139.5 114.1 120.3 73.3 3174.1 78.5
TWIST-Real-Smo 2085 1557 1874 117.1 119.8 96.9 97.9 80.0 2811.3 70.6
Any2Track-Real-Smo 2978 2785 2900 102.6 139.4 44.7 69.1 60.0 1978.8 82.6
GMT-Real-Dyn 2881 2530 2741 88.8 102.6 50.8 59.7 80.0 4098.6 126.3
TWIST-Real-Dyn 1501 1140 1357 108.5 142.7 71.0 95.4 33.3 4024.1 126.0
Any2Track-Real-Dyn 2753 2251 2552 114.2 128.2 40.9 49.5 80.0 3783.2 127.6
GMT-Sim-Smo - - - 121.0 137.4 105.9 118.5 20.0 2146.2 58.7
TWIST-Sim-Smo - - - 116.7 125.2 106.5 108.4 80.0 1684.6 49.4
Any2Track-Sim-Smo - - - 91.4 106.9 65.2 69.2 80.0 2115.1 88.1
GMT-Sim-Dyn - - - 87.8 137.0 74.7 114.0 0.0 3694.6 110.5
TWIST-Sim-Dyn - - - 100.9 127.8 62.2 74.7 40.0 3438.9 110.1
Any2Track-Sim-Dyn - - - 90.9 105.7 65.6 65.5 80.0 3949.8 130.1
Table 1: Benchmark results across controllers, input types, and simulation settings. We evaluate three representative humanoid controllers (GMT, TWIST, and Any2Track) under smoothed (Smo) and dynamic (Dyn) input motions, as well as their simulation counterparts. Higher Just Dance Score (JDS) indicates better task performance, lower MPJPE and Dynamic Time Warping (DTW) indicate more accurate motion tracking and temporal alignment, and higher Success Rate (SR) reflects improved stability. Smoothness is measured using acceleration and jerk.